# Directed Homophily-Aware Graph Neural Network

### v2

@@ 1 Introduction
Heterophily refers to the case where neighboring nodes have dissimilar features or labels [34, 15]. The node-wise homophily ratio [19] quantifies the proportion of a node’s neighbors sharing the same label, which is commonly used to measure heterophily. GNNs often perform poorly on graphs with low average homophily [33], partly due to over-smoothing [13], where deeper models produce indistinguishable node representations. Several methods have been proposed to address heterophilic settings and mitigate over-smoothing [33, 31, 25, 8, 17, 24]. However, these approaches either fail to capture fluctuations in homophily across different hops [33, 24] (illustrated in Figure 1(a)) or lack the ability to make personalized, node-specific decisions [25, 31, 8, 17], limiting their effectiveness in complex neighborhood structures.
Edge directionality is another important source of information that GNNs fail to fully exploit. While the traditional approach of converting directed graphs to undirected ones [12, 28, 7] preserves intra-class relationships, it imposes artificial symmetry that can obscure important directional, inter-class information. Both forward (original) and backward (reversed) directions can contain complementary and informative signals that are critical to understand complex graph structures, such as inter-domain connections in citation networks. Encoding the two directions separately demonstrates effectiveness [27, 21, 9]. Nonetheless, these approaches still face limitations. They ignore the homophily and structure differences between neighborhoods in opposite directions, as shown in Figure 1(b). Moreover, their reliance on layer-wise fusion restricts the model’s ability to capture long-range dependencies across the graph [21, 9].
((a)) The homophily fluctuation in Chameleon.
((b)) The directional homophily difference in Citeseer-Full.
Figure 1: The average node-wise homophily ratio in different hops. Forward denotes the direction in which the graph is constructed while backward denotes the reverse direction. The line chart reveals (a) the homophily ratio does not consistently decrease with increasing distance, and (b) the homophily levels of forward and backward neighbors can be different.
@@ 2 Related Work
Directed Graph Neural Networks. Traditional GNNs often convert directed graphs into undirected ones, potentially disrupting important inter-class links (see Appendix A for an example). To address this limitation, various methods have been developed from both spectral and spatial perspectives to explicitly model directionality in graph learning. Spectral approaches include DiGCN [26], which approximates the Laplacian using personalized PageRank transition probabilities, and MagNet [32], which introduces a Magnetic Laplacian with complex-valued phase matrices to encode directionality. Building on this, Geisler et al. [4] use both the Magnetic Laplacian and directional random walk encodings as positional encodings for Transformers, while Huang et al. [10] further enhance the Magnetic Laplacian-based encoding with additional structural factors. However, these methods often apply the same transformation to both edge directions, overlooking directional differences in neighborhood homophily.
@@ 4.2 Homophily-aware gated message passing
Therefore, to capture fluctuations in local homophily and gaps between forward and backward directions simultaneously, we introduce a homophily-aware gated message-passing mechanism, as defined in Eq. (2), while the gating strategy at the combination step is inspired by Ordered GNN [24]. In each aggregation layer, a separate gating vector is updated with a homophily-aware resettable (HAR) combination module alongside the node embeddings (shown in Figure 2b), modulating the influence of forward messages from the neighborhood in a node-specific manner.
$\small\begin{split}\mathbf{m}_{v}^{(l)}&=\text{mean}\left(\left\{\mathbf{h}_{u}^{(l-1)}|u\in\mathcal{N}(v)\right\}\right),\\ \mathbf{\hat{g}}_{v}^{(l)}&=\text{cumax}_{\leftarrow}\left(\mathbf{W}^{(l)}\left[\mathbf{h}_{v}^{(l-1)}||\mathbf{m}_{v}^{(l)}\right]+\mathbf{b}^{(l)}\right),\\ \mathbf{\tilde{g}}_{v}^{(l)}&=\text{ResetGate}(\mathbf{\tilde{g}}_{v}^{(l-1)},\mathbf{\hat{g}}_{v}^{(l)}),\\ \mathbf{h}_{v}^{(l)}&=\mathbf{\tilde{g}}_{v}^{(l)}\circ\mathbf{h}_{v}^{(l-1)}+(1-\mathbf{\tilde{g}}_{v}^{(l)})\circ\mathbf{m}_{v}^{(l)},\end{split}$ (2)
where $\text{cumax}_{\leftarrow}(\cdot)=\text{cumsum}_{\leftarrow}(\text{softmax}(\cdot))$ denotes a composite operation that applies the softmax function followed by a cumulative sum. The subscript $\leftarrow$ indicates that the accumulation proceeds in a reverse (right-to-left) direction. $\mathbf{W}^{(l)}$ and $\mathbf{b}^{(l)}$ are the learnable weight and bias term in the multilayer perceptron (MLP) layer, ||||| | stands for concatenation, and $\circ$ denotes element-wise multiplication. The first equation represents the aggregation step in the message-passing layer, while the remaining three equations define the combination step.
@@
$\small\begin{split}\mathbf{\tilde{g}}_{v}^{(l)}&=\text{ResetGate}(\mathbf{\tilde{g}}_{v}^{(l-1)},\mathbf{\hat{g}}_{v}^{(l)})=\left(1-\mathbf{r}_{v}^{(l)}\right)\circ\mathbf{\tilde{g}}_{v}^{(l-1)}+\mathbf{r}_{v}^{(l)}\circ\mathbf{\hat{g}}_{v}^{(l)},\\ \mathbf{r}_{v}^{(l)}&=\text{MLP}(\mathbf{W}_{r}\mathbf{\hat{g}}_{v}^{(l)}+\mathbf{U}_{r}\mathbf{\tilde{g}}_{v}^{(l-1)}+\mathbf{b}_{r}),\end{split}$ (3)
where $\mathbf{W}_{r}$, $\mathbf{U}_{r}$ and $\mathbf{b}_{r}$ are the two learnable weights and bias term in the MLP, respectively. The values in the gating vector tend to increase when the forward neighborhood messages deviate from the current node embedding. These values are constrained within the range $[0,1]$, where higher values correspond to a reduced influence from neighborhood messages at the current layer. This selective suppression of information from heterophilic neighbors helps mitigate the over-smoothing issue commonly observed in heterophilic graphs.
To accommodate fluctuations in neighborhood homophily across different hop distances, the gating vector values are not constrained to follow a strictly increasing or decreasing trend across layers. This behavior is enabled by the resettable gating mechanism, which allows gating values to increase when the homophily ratio we infer from the embedding similarity decreases or when the incoming message is uninformative, and decrease otherwise. Such flexibility empowers the model to suppress information aggregation from all nodes in the hops with low homophily ratios, while still enabling the integration of informative signals from more distant nodes when necessary.
@@ 4.3 Structure-aware noise-tolerant fusion
$\small\begin{split}\mathbf{h}_{v}^{\text{fused}}=\sigma\left(q_{v,\leftarrow}\mathbf{h}^{(L)}_{v,\leftarrow}+q_{v,\rightarrow}\mathbf{h}^{(L)}_{v,\rightarrow}\right),\end{split}$ (4)
@@
$\small\begin{split}q_{v,\leftarrow},q_{v,\rightarrow}&=\text{Softmax}\left(\text{NTB}(\mathbf{\tilde{h}}_{v,\leftarrow}),\text{NTB}(\mathbf{\tilde{h}}_{v,\rightarrow})\right),\\ \mathbf{\tilde{h}}_{v,\leftarrow}&=\mathbf{h}_{v,\leftarrow}+\beta\text{MLP}(\mathbf{a}_{v,\leftarrow}),\\ \text{NTB}(\mathbf{\tilde{h}}_{v,\leftarrow})&=\mathbf{\tilde{h}}_{v,\leftarrow}\mathbf{W}_{1,\leftarrow}+\epsilon\cdot\text{Softplus}(\mathbf{\tilde{h}}_{v,\leftarrow}\mathbf{W}_{2,\leftarrow}),\end{split}$ (5)
where $\text{NTB}(\cdot)$ stands for the noise-tolerant block, $\mathbf{a}_{v}$ is the corresponding vector in $\mathbf{A}$ containing all outgoing neighbors of $v$, and $\epsilon\sim N(0,1)$ denotes standard Gaussian noise. $\mathbf{W}_{1,\leftarrow},\mathbf{W}_{2,\leftarrow}\in\mathbb{R}^{p\times 1}$ are learnable weights that control clean and noisy scores, respectively. Since the adjacency matrix of a directed graph is asymmetric, it can provide richer structural information than its undirected counterpart. To fully leverage this asymmetric structure in guiding the fusion process, we incorporate an adjacency-based embedding, which has been proven effective as a structural embedding in LINKX [16], into the input of $\text{NTB}(\cdot)$. This integration is governed by the hyperparameter $\beta$.
@@
$\small\begin{split}\text{Importance}(\mathbf{h}_{\leftarrow})&=\sum_{v\in\mathcal{V}}q_{v,\leftarrow},\\ \text{Importance}(\mathbf{h}_{\rightarrow})&=\sum_{v\in\mathcal{V}}q_{v,\rightarrow},\\ L_{imp}=\text{CV}(\text{Importance}(\mathbf{h}_{\leftarrow})&,\text{Importance}(\mathbf{h}_{\rightarrow}))^{2},\end{split}$ (6)
@@ 4.4 Objective function
The first auxiliary component is the importance loss $L_{imp}$, as described in the previous section. The second auxiliary component is the branch classification loss $L_{branch}$, which is calculated as the average of classification losses from each directional branch. Formally, $L_{branch}=(L_{branch,\leftarrow}+L_{branch,\rightarrow})/2$. We adopt cross-entropy loss for both directions.
To prevent the fusion mechanism from interfering with the individual optimization of each directional encoder, we apply stop-gradient operations to the directional embeddings after computing their respective losses and before they are fused into the final output. The two auxiliary losses are weighted by hyperparameters $\lambda_{1}$ and $\lambda_{2}$, respectively. Thus, the overall objective function is formulated as:
$\small\begin{split}L_{total}=(1-\lambda_{2})(L_{cls}+\lambda_{1}L_{imp})+\lambda_{2}L_{branch}.\end{split}$ (7)
5 Experimental Results
@@ 5.1 Datasets
We evaluate the performance of our model on two homophilic datasets and three heterophilic datasets, with varying graph densities. The homophilic datasets include Cora-ML and Citeseer-Full [2], while the heterophilic datasets comprise Chameleon, Squirrel [19], and Roman-Empire [20]. Additional dataset details are provided in the Appendix B.
5.2 Experimental setup
For those baselines with results reported in [9], we adopt the results in the paper. For DHGNN, we select the best hyperparameters from 100 runs using the Optuna hyperparameter optimization framework [1]. The selected hyperparameters are detailed in Appendix C.
5.3 Node classification
Datasets Cora-ML Citeseer-Full Chameleon Squirrel Roman-Empire
MLP 77.48$\pm$1.23 80.01$\pm$1.23 46.36$\pm$2.57 34.14$\pm$1.94 65.76$\pm$0.42
GCN 52.37$\pm$1.77 54.65$\pm$1.22 64.82$\pm$2.24 53.43$\pm$2.01 73.69$\pm$0.74
GAT 54.12$\pm$1.56 55.15$\pm$1.31 45.56$\pm$3.16 39.14$\pm$2.88 71.16$\pm$0.63
GraphSAGE 54.12$\pm$1.56 55.15$\pm$1.31 45.56$\pm$3.16 39.14$\pm$2.88 71.16$\pm$0.63
H2GCN 62.86$\pm$1.45 68.34$\pm$1.36 59.39$\pm$0.98 37.90$\pm$0.02 60.11$\pm$0.52
GPRGNN 68.88$\pm$1.66 70.12$\pm$1.24 62.85$\pm$2.90 54.35$\pm$0.87 64.85$\pm$0.27
ACM-GCN 69.96$\pm$1.53 73.61$\pm$1.32 74.76$\pm$2.20 67.40$\pm$2.21 69.66$\pm$0.62
GloGNN 73.78$\pm$1.69 76.13$\pm$1.14 57.88$\pm$1.76 71.21$\pm$1.84 59.63$\pm$0.69
LINKX 72.32$\pm$1.41 78.75$\pm$1.34 68.42$\pm$1.32 61.81$\pm$1.80 37.55$\pm$0.36
FSGNN 67.51$\pm$1.65 66.35$\pm$1.16 78.27$\pm$1.28 74.10$\pm$1.89 79.92$\pm$0.56
$G^{2}$ 76.63$\pm$1.63 80.36$\pm$1.18 71.40$\pm$2.38 64.26$\pm$2.38 82.16$\pm$0.78
Di-GCN 87.36$\pm$1.06 92.75$\pm$0.75 52.24$\pm$3.65 37.74$\pm$1.54 52.71$\pm$0.32
MagNet 85.26$\pm$1.05 93.38$\pm$0.89 58.32$\pm$2.87 39.01$\pm$1.93 88.07$\pm$0.27
Dir-GNN 84.45$\pm$1.69 92.79$\pm$0.59 79.71$\pm$1.26 75.31$\pm$1.92 91.23$\pm$0.32
NDDGNN 88.14$\pm$1.28 94.17$\pm$0.58 79.79$\pm$1.04 75.38$\pm$1.95 91.76$\pm$0.27
DHGNN (ours) 88.39$\pm$0.98 94.60$\pm$0.52 80.11$\pm$1.73 76.84$\pm$1.61 89.05$\pm$0.53
Datasets Cora-ML Citeseer-Full Chameleon Squirrel
GCN 76.24$\pm$4.25 71.16$\pm$4.41 86.03$\pm$1.53 90.64$\pm$0.49
GAT 64.58$\pm$12.28 67.28$\pm$2.90 85.12$\pm$1.57 90.37$\pm$0.45
MagNet 82.20$\pm$3.36 85.89$\pm$4.61 86.28$\pm$1.23 90.13$\pm$0.53
Dir-GNN 82.23$\pm$4.01 70.62$\pm$6.56 82.76$\pm$2.72 89.17$\pm$0.52
NDDGNN 83.04$\pm$9.56 87.93$\pm$3.92 90.30$\pm$1.82 92.40$\pm$1.01
DHGNN (ours) 98.11$\pm$1.84 89.70$\pm$6.76 99.09$\pm$2.02 96.21$\pm$3.42
We evaluate our model on five datasets using the same data splits as in [9]. The results, presented in Table 1, report the mean classification accuracy on the test nodes over 10 random splits.
5.4 Link prediction
In addition to node classification, we assess the performance of the proposed DHGNN on link prediction tasks. Results are reported in Table 2. Our model achieves the highest prediction accuracy across all datasets, with absolute improvements of 15.07%, 1.77%, 8.79%, and 3.81% over the best-performing baseline on each dataset, respectively. These results indicate that the proposed method effectively captures structural information from directed neighborhoods. This may be attributed to the fact that DHGNN maintains independent encoders for each direction, rather than performing fusion after every message passing layer in NDDGNN. This design enables each encoder to capture long-range dependencies along extended paths more effectively, thereby providing richer structural information for link prediction.
5.5 Ablation study
Table 3: Ablation study for node classification. The best result (accuracy) is highlighted in bold. +ResGate denotes the resettable homophily-aware gating mechanism for message passing, and +Fusion denotes the structure-aware fusion module.
+ResGate +Fusion +$L_{branch}$ +$L_{imp}$ Chameleon Squirrel Roman-Empire
- - - - 79.21$\pm$1.03 74.96$\pm$1.95 43.35$\pm$0.63
- - - 79.52$\pm$1.38 75.22$\pm$2.13 85.55$\pm$0.46
- - 79.98$\pm$1.68 76.38$\pm$1.57 86.50$\pm$0.36
- 79.54$\pm$1.17 76.25$\pm$1.84 88.06$\pm$0.44
80.11$\pm$1.73 76.84$\pm$1.61 89.05$\pm$0.53
To evaluate the effectiveness of different components in our proposed framework, we conduct an ablation study focusing on four key elements: the resettable homophily-aware gating mechanism (ResGate), the structure-aware noise-tolerant fusion module (Fusion), and the two auxiliary losses, i.e., $L_{branch}$ and $L_{imp}$. When all components are disabled (the first row in Table 3), the model degenerates to a two-branch vanilla GCN that encodes each direction individually, and a linear layer followed by summation as the alternative fusion module. The results in Table 3 demonstrate that the homophily-aware gating mechanism significantly improves performance on heterophilic datasets. Additionally, the structure-aware fusion module enhances the extraction of valuable information from the directional encoders. The results also suggest that the combined use of the two auxiliary losses can lead to synergistic effects, further boosting classification accuracy.
5.6 Visualization of gating vectors
((a)) Forward/backward gating values on Chameleon.
((b)) Forward/backward gating values on Citeseer-Full.
Figure 3: The gating values of the first five layers on heterophilic and homophilic datasets.
Additionally, we observe a consistent decrease in most gating values from layer 1 to layer 2 on the Chameleon dataset. This aligns with the higher homophily ratio at the second hop, as illustrated in Figure 1(a). The larger variance in gating values, as indicated by the taller boxplots, may be attributed to the increased variability in neighborhood homophily at greater hop distances. These findings indicate that the proposed gating mechanism not only distinguishes between directional neighborhoods but also dynamically adapts to the non-monotonic variations in homophily across layers. Similar gating behavior is observed on other datasets, as visualized in Appendix D.
5.7 Alleviating Over-smoothing in DHGNN
((a)) Node classification accuracy on Chameleon at deeper layers, better visualised in color.
((b)) Gating values on Chameleon at deeper layers.
As shown in Figure 4(a), DHGNN effectively mitigates the over-smoothing problem. Node classification accuracy does not decrease, and even slightly improves, as the model goes deeper. The change in accuracy is obviously smaller than that observed in traditional GNNs, and remains lower than other baselines. Analysis of the gating vectors at deeper layers (Figure 4(b)) shows that while some gates saturate and refuse neighborhood messages, others remain open with lower values. This suggests that the model continues to capture informative signals from distant nodes, even at greater depths.
6 Conclusion
Appendix A Differences Between Directed and Undirected Graphs
((a)) Directed Chameleon dataset.
((b)) Undirected Chameleon dataset.
Figure A1: Class connection matrices of the directed and undirected versions of the same graph. The matrices tend to be asymmetric for heterophilic graphs, such as Chameleon.
Figure A1 illustrates how converting a directed graph to an undirected one can disrupt original inter-class connectivity patterns. In the heterophilic Chameleon dataset, this conversion removes the asymmetry between the edge ratios from class 1 to class 0 and vice versa, despite these ratios differing significantly in the original directed graph.
Appendix B Dataset Description
We provide detailed information on how each graph is constructed, along with their respective properties and statistics below:
Chameleon [19]: This dataset represents a Wikipedia page-page network on the topic of chameleons. Nodes correspond to Wikipedia articles, and edges denote mutual links between these articles. Node features indicate the presence of particular nouns in the articles. The nodes are classified into five categories based on their average monthly traffic. This dataset contains 2,277 nodes and 36,101 edges.
Squirrel [19]: Similar to the Chameleon dataset, the Squirrel dataset represents a Wikipedia page-page network, but on the topic of squirrels. It also consists of nodes representing Wikipedia articles and edges representing mutual links. Node features are based on the presence of informative nouns in the text of the articles. The Squirrel dataset has 5,201 nodes and 217,073 edges.
Cora-ML [2]: This dataset is a citation network where nodes represent scientific publications, and edges represent citations between them. The dataset is derived from the Cora dataset, but with multi-label classification. It contains 2,995 nodes and 8,416 edges. Each node is assigned multiple labels that represent the categories of the papers.
Citeseer-Full [2]: The Citeseer-Full dataset is another citation network where nodes represent scientific publications and edges represent citations between them. This dataset includes the entire Citeseer dataset without filtering. It contains 3,312 nodes and 4,732 edges, with each node labeled according to its category.
Roman-Empire [20]: The dataset is based on the Roman Empire article from English Wikipedia, which was selected since it is one of the longest articles on Wikipedia. Each node in the graph corresponds to one (non-unique) word in the text. Thus, the number of nodes in the graph is equal to the article’s length. Two words are connected with an edge if at least one of the following two conditions holds: either these words follow each other in the text, or these words are connected in the dependency tree of the sentence (one word depends on the other). It contains 22662 nodes and 32927 edges. The class of a node is its syntactic role (The 17 most frequent roles were selected as unique classes and group all the other roles into the 18th class). For node features, the dataset uses FastText word embeddings [6].
Table A1: Dataset properties and statistics.
@@ Appendix B Dataset Description
Appendix C Experiment Settings

### v3

@@ Abstract
keywords:
Graph Neural Networks , Heterophily , Directed Graph
journal: Information Fusion
[…]
[label1]organization=College of Computing and Data Science, Nanyang Technological University, addressline=50 Nanyang Avenue, postcode=639798, country=Singapore
[…]
[label2]organization=Institute for Infocomm Research, A*STAR, addressline=2 Fusionopolis Way, Innovis 08-01, postcode=138634, country=Singapore
footnotetext: Our code is available at: https://anonymous.4open.science/r/DHGNN-7FE5.
@@ 1 Introduction
Heterophily refers to the case where neighboring nodes have dissimilar features or labels [37, 16, 27, 32, 8]. The node-wise homophily ratio [20] quantifies the proportion of a node’s neighbors sharing the same label, which is commonly used to measure heterophily. GNNs often perform poorly on graphs with low average homophily [36], partly due to over-smoothing [15], where deeper models produce indistinguishable node representations. Several methods have been proposed to address heterophilic settings and mitigate over-smoothing [36, 34, 26, 10, 18, 25]. However, these approaches either fail to capture fluctuations in homophily across different hops [36, 25] (illustrated in Figure 1(a)) or lack the ability to make personalized, node-specific decisions [26, 34, 10, 18], limiting their effectiveness in complex neighborhood structures.
(a) The homophily fluctuation in Chameleon.
(b) The directional homophily difference in Citeseer-Full.
Figure 1: The average node-wise homophily ratio in different hops. Forward denotes the direction in which the graph is constructed while backward denotes the reverse direction. The line chart reveals (a) the homophily ratio does not consistently decrease with increasing distance, and (b) the homophily levels of forward and backward neighbors can be different. We further explore this phenomenon with an example in A.
(a) Directed Chameleon dataset.
(b) Undirected Chameleon dataset.
Figure 2: Class connection matrices of the directed and undirected versions of the same graph. The matrices tend to be asymmetric for heterophilic graphs, such as Chameleon. Converting this graph from directed to undirected could eliminate the asymmetry between the edge ratios from class 1 to class 0 and from class 0 to class 1, while the two ratios differ significantly in the original directed graph.
Edge directionality is another important source of information that GNNs fail to fully exploit. While the traditional approach of converting directed graphs to undirected ones [14, 30, 9] preserves intra-class relationships, it imposes artificial symmetry that can obscure important directional, inter-class information, as illustrated in Figure 2. Both forward (original) and backward (reversed) directions can contain complementary and informative signals that are critical to understand complex graph structures, such as inter-domain connections in citation networks. Encoding the two directions separately demonstrates effectiveness [29, 22, 11]. Nonetheless, these approaches still face limitations. They ignore the homophily and structure differences between neighborhoods in opposite directions, as shown in Figure 1(b). Moreover, their reliance on layer-wise fusion restricts the model’s ability to capture long-range dependencies across the graph [22, 11].
@@ 2 Related Work
Directed Graph Neural Networks. Traditional GNNs often convert directed graphs into undirected ones, potentially disrupting important inter-class links. To address this limitation, various methods have been developed from both spectral and spatial perspectives to explicitly model directionality in graph learning. Spectral approaches include DiGCN [28], which approximates the Laplacian using personalized PageRank transition probabilities, and MagNet [35], which introduces a Magnetic Laplacian with complex-valued phase matrices to encode directionality. Building on this, [5] uses both the Magnetic Laplacian and directional random walk encodings as positional encodings for Transformers, while [12] further enhances the Magnetic Laplacian-based encoding with additional structural factors. However, these methods often apply the same transformation to both edge directions, overlooking directional differences in neighborhood homophily.
@@ 3 Preliminaries
Homophily. Homophily is the tendency of connected nodes in a graph to share similar labels or features, and it underpins the success of many GNNs. A common way to measure this is the node-wise homophily ratio, defined for a node $v$ with label $y_{v}$ as
$h_{v}=\frac{\left|\left\{u\in\mathcal{N}(v)\;\middle|\;y_{u}=y_{v}\right\}\right|}{\left|\mathcal{N}(v)\right|},$ (1)
where $\mathcal{N}(v)$ is the set of neighbors of $v$. The overall graph homophily is computed by averaging $h_{v}$ over all nodes. High homophily typically benefits standard GNNs, while low homophily (high heterophily) poses challenges for conventional message passing.
In scenarios where node labels are partially or entirely unavailable, such as in link prediction tasks, the notion of homophily can be extended to quantify feature similarity between neighboring nodes.
@@ 4.2 Homophily-aware gated message passing
Therefore, to capture fluctuations in local homophily and gaps between forward and backward directions simultaneously, we introduce a homophily-aware gated message-passing mechanism, as defined in Eq. (3). In each aggregation layer, a separate gating vector is updated with a homophily-aware resettable (HAR) combination module alongside the node embeddings (shown in Figure 3b), modulating the influence of forward messages from the neighborhood in a node-specific manner.
$\begin{split}\mathbf{m}_{v}^{(l)}&=\text{mean}\left(\left\{\mathbf{h}_{u}^{(l-1)}|u\in\mathcal{N}(v)\right\}\right),\\ \mathbf{\hat{g}}_{v}^{(l)}&=\text{cumax}_{\rightarrow}\left(\mathbf{W}^{(l)}\left[\mathbf{h}_{v}^{(l-1)}||\mathbf{m}_{v}^{(l)}\right]+\mathbf{b}^{(l)}\right),\\ \mathbf{\tilde{g}}_{v}^{(l)}&=\text{ResetGate}(\mathbf{\tilde{g}}_{v}^{(l-1)},\mathbf{\hat{g}}_{v}^{(l)}),\\ \mathbf{h}_{v}^{(l)}&=\mathbf{\tilde{g}}_{v}^{(l)}\circ\mathbf{h}_{v}^{(l-1)}+(1-\mathbf{\tilde{g}}_{v}^{(l)})\circ\mathbf{m}_{v}^{(l)},\end{split}$ (3)
where $\text{cumax}_{\rightarrow}(\cdot)=\text{cumsum}_{\rightarrow}(\text{softmax}(\cdot))$ denotes a composite operation that applies the softmax function followed by a cumulative sum. The subscript $\rightarrow$ indicates that the accumulation proceeds in a left-to-right direction. $\mathbf{W}^{(l)}$ and $\mathbf{b}^{(l)}$ are the learnable weight and bias term in the multilayer perceptron (MLP) layer, $||$ stands for concatenation, and $\circ$ denotes element-wise multiplication. The first equation represents the aggregation step in the message-passing layer, while the remaining three equations define the combination step.
@@
$\begin{split}\mathbf{\tilde{g}}_{v}^{(l)}&=\text{ResetGate}(\mathbf{\tilde{g}}_{v}^{(l-1)},\mathbf{\hat{g}}_{v}^{(l)})=\left(1-\mathbf{r}_{v}^{(l)}\right)\circ\mathbf{\tilde{g}}_{v}^{(l-1)}+\mathbf{r}_{v}^{(l)}\circ\mathbf{\hat{g}}_{v}^{(l)},\\ \mathbf{r}_{v}^{(l)}&=\sigma(\mathbf{W}_{r}\mathbf{\hat{g}}_{v}^{(l)}+\mathbf{U}_{r}\mathbf{\tilde{g}}_{v}^{(l-1)}+\mathbf{b}_{r}),\end{split}$ (4)
where $\sigma(\cdot)$ stands for the sigmoid function, and $\mathbf{W}_{r}$, $\mathbf{U}_{r}$ and $\mathbf{b}_{r}$ are the learnable weights and the bias term, respectively. The values in the gating vector tend to increase when the forward neighborhood messages deviate from the current node embedding. These values are constrained within the range $[0,1]$, where higher values correspond to a reduced influence from neighborhood messages at the current layer. This selective suppression of information from heterophilic neighbors helps mitigate the over-smoothing issue commonly observed in heterophilic graphs.
While the gating strategy at the combination step is inspired by Ordered GNN [25], the original design assumes that the neighborhood homophily ratio strictly decreases with increasing hop distance. In contrast, our gating mechanism is designed to accommodate fluctuations in neighborhood homophily across different hops. To this end, the gating vector values are not constrained to follow a strictly monotonic trend across layers. This behavior is enabled by the resettable gating mechanism, which allows gating values to increase when the homophily ratio we infer from the embedding similarity decreases or when the incoming message is uninformative, and decrease otherwise. Such flexibility enables the model to gradually attenuate less relevant signals from distant nodes, while still retaining useful long-range information when necessary. As a result, DHGNN can dynamically adapt its receptive field depending on the directionality and structure of the graph, leading to more expressive and context-aware node representations. We provide an analysis of the monotonicity of gating values in B.
@@ 4.3 Structure-aware noise-tolerant fusion
[…]
@@
$\begin{split}q_{v,\leftarrow},q_{v,\rightarrow}&=\text{Softmax}\left(\text{NTB}(\mathbf{\tilde{h}}_{v,\leftarrow}),\text{NTB}(\mathbf{\tilde{h}}_{v,\rightarrow})\right),\\ \mathbf{\tilde{h}}_{v,\leftarrow}&=\mathbf{h}_{v,\leftarrow}+\beta\text{MLP}(\mathbf{a}_{v,\leftarrow}),\\ \text{NTB}(\mathbf{\tilde{h}}_{v,\leftarrow})&=\mathbf{\tilde{h}}_{v,\leftarrow}\mathbf{W}_{1,\leftarrow}+\epsilon\cdot\text{Softplus}(\mathbf{\tilde{h}}_{v,\leftarrow}\mathbf{W}_{2,\leftarrow}),\end{split}$ (6)
where $\text{NTB}(\cdot)$ stands for the noise-tolerant block, $\mathbf{a}_{v}$ is the corresponding vector in $\mathbf{A}$ containing all outgoing neighbors of $v$, and $\epsilon\sim N(0,1)$ denotes standard Gaussian noise. We introduce this noise term in the fusion mechanism to promote a stable and decisive combination of the two directional representations. $\mathbf{W}_{1,\leftarrow},\mathbf{W}_{2,\leftarrow}\in\mathbb{R}^{p\times 1}$ are learnable weights that control clean and noisy scores, respectively. Since the adjacency matrix of a directed graph is asymmetric, it can provide richer structural information than its undirected counterpart. To fully leverage this asymmetric structure in guiding the fusion process, we incorporate an adjacency-based embedding, which has been proven effective as a structural embedding in LINKX [17], into the input of $\text{NTB}(\cdot)$. This integration is governed by the hyperparameter $\beta$.
@@
$\begin{split}\text{Importance}(\mathbf{h}_{\leftarrow})&=\sum_{v\in\mathcal{V}}q_{v,\leftarrow},\\ \text{Importance}(\mathbf{h}_{\rightarrow})&=\sum_{v\in\mathcal{V}}q_{v,\rightarrow},\\ L_{imp}=\text{CV}(\text{Importance}(\mathbf{h}_{\leftarrow})&,\text{Importance}(\mathbf{h}_{\rightarrow}))^{2},\end{split}$ (7)
@@ 4.4 Objective function
The first auxiliary component is the importance loss $L_{imp}$, as described in the previous section. The second auxiliary component, the branch classification loss $L_{branch}$, encourages each branch to independently learn information useful for node classification. It is calculated as the average of classification losses from each directional branch. Formally, $L_{branch}=(L_{branch,\leftarrow}+L_{branch,\rightarrow})/2$. We adopt cross-entropy loss for both directions.
To prevent the fusion mechanism from interfering with the individual optimization of each directional encoder, we apply stop-gradient operations to the directional embeddings after computing their respective losses and before they are fused into the final output. The stop-gradient operation effectively prevents gradients from propagating back through the encoder for each branch, ensuring that the previous quantities (such as the node embeddings from each encoder) are treated as fixed targets rather than learnable inputs during the update step. The two auxiliary losses are weighted by hyperparameters $\lambda_{1}$ and $\lambda_{2}$, respectively. $\lambda_{2}$ is introduced to form the total loss as a convex combination of the losses from the fused output and the individual branches. Thus, the overall objective function is formulated as:
$\begin{split}L_{total}=(1-\lambda_{2})(L_{cls}+\lambda_{1}L_{imp})+\lambda_{2}L_{branch}.\end{split}$ (8)
5 Experiments
@@ 5.1 Datasets
We evaluate our model on five representative benchmarks. Cora-ML and Citeseer-Full [2] are homophilic citation networks whose nodes represent scientific publications with category annotations and edges denote citation links. Chameleon and Squirrel [20] are heterophilic Wikipedia page–page networks where nodes are articles connected by mutual links, equipped with bag-of-nouns features and traffic-based labels. Roman-Empire [21] is a word-level heterophilic graph constructed from a long Wikipedia article, where nodes are tokens linked by textual adjacency or dependency relations, with FastText-based features [7] and syntactic-role labels. Additional dataset details are provided in Table 1.
Table 1: Dataset properties and statistics.
@@
5.2 Experimental setup
For those baselines with results reported in [11], we adopt the results in the paper. For DHGNN, we select the best hyperparameters from 100 runs using the Optuna hyperparameter optimization framework [1]. The selected hyperparameters are detailed in C.
We conduct our experiments on an Intel Core i9-10940X CPU and 2 NVIDIA RTX A5000 GPUs.
5.3 Node classification
Datasets Cora-ML Citeseer-Full Chameleon Squirrel Roman-Empire
[…]
Datasets Cora-ML Citeseer-Full Chameleon Squirrel
[…]